# LiSA: Leveraging Link Recommender to Attack Graph Neural Networks via Subgraph Injection

### v1

@@ Abstract.
Graph Neural Network, Adversarial Attack, Subgraph Injection Attack
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Security and privacyccs: Networks
1. Introduction
In recent years, the field of machine learning-based artificial intelligence has experienced exponential growth across various dynamic domains, including computer vision and natural language processing. Graph Neural Networks (GNNs) (Zhou et al., 2018), as a powerful deep learning algorithm, have shown remarkable proficiency in modeling graph-structured data. They have become indispensable in diverse applications such as social and traffic network analysis (Ying et al., 2018; Wu et al., 2020), applied sciences (Shlomi et al., 2020; Jiang et al., 2021), and financial systems (Wang et al., 2019). GNNs primarily operate on a message-passing mechanism that aggregates and integrates information from neighboring nodes into the target nodes (Xu et al., 2018). This process of iterative information aggregation enables GNNs to effectively update node representations while incorporating the structural properties. Additionally, through the computation of node embeddings as an encoder and the careful construction of a decoder, GNNs can effectively approximate the adjacency matrix of a graph, solidifying their importance as a vital element in link prediction tasks (Zhang and Chen, 2018).
@@ 1. Introduction
On the flip side, the powerful message-passing mechanism of GNNs also brings its share of challenges and limitations. Malicious attackers can intentionally perturb the graph structure and features. Such deliberately tampered information can propagate among nodes, leading to a degraded performance of GNNs (Dai et al., 2018; Zügner et al., 2018; Zügner and Günnemann, 2019; Chen et al., 2018; Wang et al., 2020; Chang et al., 2019). Recognizing the importance of designing trustworthy GNNs, investigating various attack methods has become an essential initial undertaking (Dai et al., 2022b). Consequently, there is a growing focus on adversarial attacks targeting GNNs within academic and industrial communities. Among these, manipulation attacks (Zügner et al., 2018) have been demonstrated to be particularly effective. These attacks involve adding or deleting edges, or modifying node features in the original graph, all within a constrained budget.
@@
Nonetheless, adversaries frequently encounter practical hurdles when attempting to manipulate these data, as it usually requires direct access to the backend data, a task that is not easily achievable. On the other hand, node injection attacks (NIAs) (Sun et al., 2020; Tao et al., 2021; Zou et al., 2021) present a more feasible approach. Such attacks degrade classification performance by introducing malicious nodes linked to target nodes, spreading harmful attributes. Within citation networks, adversaries may craft counterfeit papers and insert references to target papers, ultimately causing them to be misclassified. This kind of emerging attack method have demonstrated high success rates and are increasingly seen as a more practical method in real-world scenarios. Injecting fake nodes is often more feasible than modifying data on existing nodes, making these techniques particularly insidious.
@@
There is no doubt about the practicality and remarkable effectiveness of NIAs. Nevertheless, current research often focuses on ideal conditions, which implies that adversaries can freely establish connections between target nodes and injected fake nodes. However, this assumption does not universally hold across various applications. Especially in applications where security is critical, like social networks, fake users who connect with the target user can subtly alter the ground truth label of the target user, leading to targeted advertising. However, it is not certain that target users will approve connection request with fake accounts. Thus, evaluating these attack models under the assumption of cost-free linking can be misleading, potentially leading to an overestimation of the attack outcomes. These concerns highlighted above have motivated us to focus on enhancing the likelihood of adversaries successfully linking fake nodes to the target nodes. Given the prevalence of link recommendation algorithms within GNN-based applications, we explore the potential to exploit these link recommenders to facilitate malicious link formation. The proposed subgraph injection attack (SIA) scenario is shown in Fig. 1. Our method does not manually add edges between target victim nodes and the aggressor fake nodes. Instead, it deceives the link recommender of GNN into actively initiating links formation. Once these links are established, the classification performance on the target victim nodes suffers a considerable decline, ultimately enabling a successful attack. This novel approach to adversarial attacks is crucial for real-world applications yet remains largely unexplored. Moreover, the uncertain feasibility of attacking multiple GNN models simultaneously poses a significant challenge in solving this problem.
Figure 1. SIA scenario via link recommender.
@@
In this work, we investigate a unique SIA scenario, which injects an isolated subgraph to deceive both the link recommender and the node classifier in a GNN simultaneously. Specifically, it aims to mislead the link recommender into creating adversarial links from targeted victim nodes to the subgraph, thereby undermining the accuracy of node classification. Our contributions in this study are outlined as follows:
  • We explore an innovative attack scenario utilizing the link recommender to compromise the node classification accuracy within a GNN.
  • We introduce Link Recommender-Subgraph Injection Attack (LiSA), a pioneering framework crafted to effectively generate subgraphs by employing a dual surrogate model and bi-level optimization.
  • Through rigorous experimentation on diverse real-world datasets, we validate the effectiveness and applicability of our proposed methodologies.
2. Related Works
2.1. Graph Neural Networks
GNNs have emerged as a powerful tool for extracting insights from graph-structured data, serving crucial roles in tasks such as node classification and link prediction. Among these, the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) stands out for node classification tasks. GCN employs a localized first-order approximation of spectral graph convolutions, facilitating the efficient learning of node representations that capture the underlying structure of the graph. Simplifying this further, Simple Graph Convolution (SGC) (Wu et al., 2019) reduces complexity and parameter count by merging convolution and linear transformation into a single operation. Additionally, GraphSAGE (Hamilton et al., 2017) has become notable for its capability to generate node embeddings by sampling and aggregating features from the local neighborhood of a node, demonstrating proficiency in handling large-scale graphs. In the domain of link prediction, both Graph Autoencoder (GAE) and its variant, Variational Graph Autoencoder (VGAE), have shown considerable efficacy (Kipf and Welling, 2016). GAE focuses on learning low-dimensional node representations to reconstruct adjacency matrix of a graph, thereby predicting missing links. VGAE expands on this by integrating a variational autoencoder framework, introducing a probabilistic modeling of latent variable distributions. This enhances the accuracy and robustness of link prediction models by capturing the inherent uncertainty in graph structures.
2.2. Adversarial Attacks on Graph Neural Networks
Adversarial attacks on GNNs aim to exploit model vulnerabilities, resulting in errors or inaccurate predictions. These attacks are generally categorized into evasion, poisoning, and backdoor attacks, each characterized by distinct methodologies and consequences.
@@ 2.2. Adversarial Attacks on Graph Neural Networks
Evasion attacks occur during the test phase, where adversaries manipulate the input graph data to lead the GNN to make incorrect predictions (Zügner et al., 2018; Zou et al., 2021; Xu et al., 2019). These manipulations could involve strategic alterations in the graph structure or node features, significantly affecting the performance of the GNN. The stealth of evasion attacks, which do not require changes to the trained model, makes them particularly challenging to detect. Following evasion attacks, poisoning attacks are conducted during the training phase of the GNN. In this scenario, attackers inject the training set with crafted data (Dai et al., 2018; Sun et al., 2020), either by adding, deleting edges, or modifying nodes, thereby maliciously altering the data. These alterations subtly distort the training process, leading the model to unwittingly integrate these adversarial modifications. The performance of a model consequently deteriorates, introducing specific errors or biases, undermining its reliability and accuracy. Lastly, backdoor attacks are implemented by inserting hidden triggers in the GNN during its training phase (Xi et al., 2020; Zhang et al., 2021; Dai et al., 2023). When activated in new data, these triggers cause the model to generate incorrect results, while maintaining normal functionality under standard conditions. This makes backdoor attacks particularly deceptive.
@@
However, none of the existing attack methods can fully demonstrate their capabilities in practical scenarios. Attackers do not have the capability to modify the original graph in evasion and poisoning attacks. Even with the utilization of relatively accessible malicious injection techniques in these attack types, the failure to generate links immediately undermines the attack, leading to a considerable degradation in their efficacy. Therefore, we propose leveraging another fundamental component of a GNN system: the link recommender. By incorporating the link recommender, we aim to facilitate the generation of links between the target victim and injected malicious nodes. This integration enhances the likelihood of establishing connections between the aggressor subgraph and the target victim nodes, consequently leading to a higher success rate of practical attacks.
3. Preliminaries
@@
In this section, we begin by introducing the notations used in this paper. Subsequently, we lay the groundwork by outlining the fundamental concepts of node classification and link prediction tasks in GNNs. Building upon this foundation, we then elaborate on the threat model and problem formulation related to the specific attack scenario of the SIA, thereby establishing the basis for our subsequent analysis. For preliminary analysis, we investigate a heuristic technique known as GraphCopy to assess the viability of the proposed SIA scenario and establish a foundational benchmark.
3.1. Notations
@@
(1) $A_{p}=\left[\begin{array}[]{cc}A_{o}&0\\ 0&A_{s}\\ \end{array}\right],\quad F_{p}=\left[\begin{array}[]{cc}F_{o}\\ F_{s}\\ \end{array}\right]$
@@
3.2. GNN for Node Classification
The objective of node classification in GNNs is to predict the category of a node using its features and the surrounding graph structure. Consider a graph $G=(V,E)$, where $V$ and $E$ represent nodes and edges, respectively. Each node $v\in V$ is associated with a feature vector $\mathbf{x}_{v}$. The nodes aggregate information from their neighbors, expressed as:
(2) $\mathbf{{h}}_{v}^{(l+1)}={AGGREGATE}^{(l)}(\{\mathbf{{h}}_{u}^{(l)}:u\in\mathcal{N}(v)\})$
Here, $\mathbf{{h}}_{v}^{(l)}$ denotes the feature representation of node $v$ at layer $l$, and $\mathcal{N}(v)$ represents the neighbors of $v$. The aggregated features undergo a transformation through a neural network layer, and the outputs of the final layer are utilized for node classification:
(3) $\hat{y_{v}}=f(\mathbf{{h}}_{v}^{(L)})$
where $f$ denotes the activation function, $\hat{y_{v}}$ the predicted label for node $v$, and $L$ the final layer.
3.3. GNN for Link Prediction
Link prediction in GNNs focuses on assessing the probability of an edge existing between two nodes in a graph. Initially, the model generates feature representations for each node, then a linking score is decoded for each possible edge and consequently transformed into probabilities using a sigmoid function:
(4) $p_{uv}=\sigma(\mathbf{z}_{u}^{T}\cdot\mathbf{z}_{v})$
where $\mathbf{z}_{u}$ and $\mathbf{z}_{v}$ denote the representations of node $u$ and $v$ obtained by the GNN encoder, and $p_{uv}$ represents the probability of an edge existing between nodes $u$ and $v$.
3.4. Threat Model and Problem Formulation
@@
Problem. We assume the presence of a link recommender $f_{\phi}$ in the GNN, utilizing a subset of edges from core users in the original graph to suggest connections, relying on the highest linking score as determined by Eq. 4. This results in an updated graph $G_{r}$, with:
(5) $G_{r}=(\hat{A},F_{p}),\quad\hat{A}=f_{\phi}(G_{p})$
@@
(6) $\hat{A}=\left[\begin{array}[]{cc}A_{o}&A_{l}\\ {A}_{l}^{T}&A_{s}\\ \end{array}\right]$
@@ 3.5. Preliminary Analysis
(8) $G_{s}=\text{Copy}(\mathcal{N}_{n}(v_{t})),\quad\mathcal{N}_{n}(v_{t})=\{u\in V|\text{dist}(u,v_{t})\leq n\}$
Here, $\text{Copy}()$ is the replicating operation, $\mathcal{N}_{n}(v_{t})$ represent the n-hop neighborhoods set on $v_{t}$, and $\text{dist}(u,v_{t})$ gives the minimum distance between nodes $u$ and $v_{t}$. Such similarity seemingly facilitates the formation of links between the original graph and subgraph. However, simply generating links does not achieve our primary goal of reducing the accuracy of node classification. To address this, we introduce a perturbation in the form of Gaussian noise to the node features of the duplicated subgraph:
(9) $\widetilde{F_{s}}=F_{s}+\mathcal{N}(\mu,\sigma^{2})$
Here, $\widetilde{F_{s}}$ represents the duplicated subgraph features after a Gaussian perturbation $\mathcal{N}(\mu,\sigma^{2})$, which aims to hinder the classification accuracy. By fine-tuning the number of hops and the intensity of perturbation, we can achieve a balance that facilitates link generation while reducing node classification accuracy, thereby achieving the adversarial objectives.
@@
GraphCopy validates the feasibility of the SIA scenario and accomplishes the predetermined adversarial objectives. However, it faces two significant limitations that reduce its effectiveness and practicality in real-world scenarios. First, simply duplicating the graph and altering node features does not ensure a high success rate for adversarial attacks. Second, replicating the n-hop neighborhood of a target node leads to a subgraph size that increases quadratically with the node degree. For nodes with high degrees, the attack budget increase substantially, making GraphCopy impractical for use in dense or large-scale networks. The results of the attack are provided in Appendix A and will be analyzed in the following sections.
4. Methodology
@@
In this section, we investigate an optimization-based approach for generating subgraphs to realize the SIA scenario. While extensive research has focused on targeting either node classification or link prediction models independently (Trappolini et al., 2024), the exploitation of the interaction between these two distinct model types to achieve an adversarial objective remains largely unexplored. Moreover, achieving dual adversarial goals within the SIA scenario presents significant technical challenges. Intuitively, generating links is more likely when there is a close resemblance in the node representations between the target victim nodes and the nodes of the aggressor subgraph, whereas a greater disparity in these representations increases the chances of the victim nodes being misclassified. As a result, a more deliberate subgraph optimization technique is essential to balance two conflicting objectives and identify the optimal subgraph for SIA. To this end, we introduce the LiSA framework, as depicted in Fig. 2.
Figure 2. An overview of LiSA framework.
4.1. LiSA Framework
@@
The LiSA framework is engineered to challenge the performance of two GNN models through a strategic deployment of dual-surrogate models for bi-level optimization in a black-box attack context. We select the GAE and GCN as the surrogate models for link prediction and node classification, aiming to refine the adjacency matrix and feature attributes of the subgraph to meet two critical optimization goals concurrently: (i) the generation of adversarial edges between the original graph and the subgraph via the link recommender, and (ii) negatively affect node classification performance on target victim nodes once they are connected to the subgraph. To navigate these objectives, the attack loss function of LiSA includes two terms, balanced by a factor $\alpha$, to regulate the contribution of each model:
(10) $\mathcal{L}_{atk}=\mathcal{L}^{cls}_{atk}(G_{r};\theta)+\alpha\mathcal{L}^{link}_{atk}(G_{p};\phi)$
@@
(11) $\mathcal{L}^{link}_{atk}=\mathcal{L}_{recon}=-\sum_{i\in V}\sum_{j\in V}\log\hat{a}_{ij}-\sum_{n=1}^{Q}\cdot\mathrm{E}_{v_{n}\sim P_{n}(v_{i})}\log(1-\hat{{a_{n}}}_{ij})$
@@
Notably, the approach begins by pretraining a node classifier using a limited label set, subsequently altering the top-2 output classes to define the attack labels (Mopuri et al., 2020). This approach circumvents the traditional use of negative cross-entropy loss used in GradMax, which complicates the task of balancing the two objectives during subgraph training. The continuously increasing adversarial loss for classification, $\mathcal{L}^{cls}_{atk}$, usually fails to maintain balance between these two loss terms. By giving precedence to the second-highest output class for the attack, we foster a more stable training environment. The attack label can also be switched to target a specific class for a label-targeted attack. The comprehensive training strategy is outlined as follows:
(13) $\min_{A_{s},F_{s}}$ $\mathcal{L}^{cls}_{atk}(G_{r};\theta^{*})+\alpha\mathcal{L}^{link}_{atk}(G_{p};\phi^{*})$
s.t. $\theta^{*}=\mathop{\arg\min}_{\theta}\mathcal{L}^{cls}_{train}(f_{\theta}(G_{r})),$
$\phi^{*}=\mathop{\arg\min}_{\phi}\mathcal{L}^{link}_{train}(f_{\phi}(G_{p})).$
@@
The method involves the simultaneous bi-level optimization of the two surrogate models, with optimizing space limited to the subgraph. It also ensures the node degree and features closely resemble those of the original graph, making the alterations difficult to detect.
4.2. Training Algorithm

### v2

@@ Abstract
Keywords:
Graph Neural Network Adversarial Attack
1 Introduction
@@
In recent years, Graph Neural Networks (GNNs) have shown remarkable proficiency in modeling graph-structured data. GNNs primarily operate on a message-passing mechanism. This process of iterative information aggregation enables GNNs to effectively update node representations while incorporating the structural properties. However, the powerful message-passing mechanism of GNNs also brings its share of challenges and limitations. Malicious attackers can intentionally perturb the graph structure and features. Such deliberately tampered information can propagate among nodes, leading to a degraded performance of GNNs [6, 20, 21, 2, 14, 1]. Recognizing the importance of designing trustworthy GNNs, investigating various attack methods has become an essential initial undertaking [5]. Consequently, there is a growing focus on adversarial attacks targeting GNNs within academic and industrial communities. Among these, manipulation attacks [20] have been demonstrated to be particularly effective. These attacks involve adding or deleting edges, or modifying node features in the original graph, all within a constrained budget.
@@
Nonetheless, adversaries frequently encounter practical hurdles when attempting to manipulate these data, as it usually requires direct access to the backend data, a task that is not easily achievable. On the other hand, node injection attacks (NIAs) [11, 12, 19] present a more feasible approach. Such attacks degrade classification performance by introducing malicious nodes linked to target nodes, spreading harmful attributes. Within citation networks, adversaries may craft counterfeit papers and insert references to target papers, ultimately causing them to be misclassified. This kind of emerging attack method have demonstrated high success rates and are increasingly seen as a more practical method in real-world scenarios. Injecting fake nodes is often more feasible than modifying data on existing nodes, making these techniques particularly insidious.
@@
There is no doubt about the practicality and effectiveness of NIAs. Nevertheless, current research often focuses on ideal conditions, which implies that adversaries can freely establish connections between target nodes and injected fake nodes. However, this assumption does not universally hold across various applications. Especially in applications where security is critical, like social networks, fake users who connect with the target user can subtly alter the ground truth label of the target user, leading to targeted advertising. However, it is not certain that target users will approve connection request with fake accounts. Thus, evaluating these attack models under the assumption of cost-free linking can be misleading, potentially leading to an overestimation of the attack outcomes. These concerns highlighted above have motivated us to focus on enhancing the likelihood of adversaries successfully linking fake nodes to the target nodes. Given the prevalence of link recommendation algorithms within GNN-based applications, we explore the potential to exploit these link recommenders to facilitate malicious link formation. The proposed subgraph injection attack (SIA) scenario is shown in Fig. 1. Our method does not manually add edges between target victim nodes and the aggressor fake nodes. Instead, it deceives the link recommender of GNN into actively initiating links formation. Once these links are established, the classification performance on the target victim nodes suffers a considerable decline, ultimately enabling a successful attack. This novel approach to adversarial attacks is crucial for real-world applications yet remains largely unexplored. Moreover, the uncertain feasibility of attacking multiple GNN models simultaneously poses a significant challenge in solving this problem.
Figure 1: SIA scenario via link recommender.
@@
In this work, we investigate a unique SIA scenario, which injects an isolated subgraph to deceive both the link recommender and the node classifier in a GNN simultaneously. Specifically, it aims to mislead the link recommender into creating adversarial links from targeted victim nodes to the subgraph, thereby undermining the accuracy of node classification. Our contributions in this study are outlined as follows: (1) We explore an innovative attack scenario utilizing the link recommender to compromise the node classification accuracy within a GNN. (2) We introduce Link Recommender-Subgraph Injection Attack (LiSA), a pioneering framework crafted to effectively generate subgraphs by employing a dual surrogate model and bi-level optimization. (3) Through rigorous experimentation on diverse real-world datasets, we validate the effectiveness and applicability of our proposed methodologies.
2 Related Works
@@
Adversarial attacks on GNNs aim to exploit model vulnerabilities, resulting in errors or inaccurate predictions. These attacks are generally categorized into evasion, poisoning, and backdoor attacks, each characterized by distinct methodologies and consequences. Evasion attacks occur during the test phase, where adversaries manipulate the input graph data to lead the GNN to make incorrect predictions [20, 19, 17]. These manipulations could involve strategic alterations in the graph structure or node features, significantly affecting the performance of the GNN. The stealth of evasion attacks, which do not require changes to the trained model, makes them particularly challenging to detect. Following evasion attacks, poisoning attacks are conducted during the training phase of the GNN. In this scenario, attackers inject the training set with crafted data [6, 11], either by adding, deleting edges, or modifying nodes. These alterations subtly distort the training process, leading the model to unwittingly integrate these adversarial modifications. The performance of a model consequently deteriorates, introducing specific errors or biases, undermining its reliability and accuracy. Lastly, backdoor attacks are implemented by inserting hidden triggers in the GNN during its training phase [16, 18, 4]. When activated in new data, these triggers cause the model to generate incorrect results, while maintaining normal functionality under standard conditions. This makes backdoor attacks particularly deceptive.
@@
Unfortunately, none of the existing attack methods can fully demonstrate their capabilities in practical scenarios. Attackers do not have the capability to modify the original graph in evasion and poisoning attacks. Even with the utilization of relatively accessible malicious injection techniques, the failure to generate links immediately undermines the attack, leading to a considerable degradation in their efficacy. Therefore, we propose leveraging the link recommender to facilitate the generation of links between the target victim and injected nodes. The involvement of link recommender enhances the likelihood of establishing connections between the aggressor subgraph and the target victim nodes, consequently leading to a higher success rate in practical attacks.
3 Preliminaries
@@
In this section, we begin by introducing the notations used in this paper. Building upon this foundation, we then elaborate on the threat model and problem formulation related to the specific attack scenario of the SIA. For preliminary analysis, we investigate a heuristic technique known as GraphCopy to assess the viability of the proposed SIA scenario and establish a foundational benchmark.
3.1 Notations
@@
$A_{p}=\left[\begin{array}[]{cc}A_{o}&0\\ 0&A_{s}\\ \end{array}\right],\quad F_{p}=\left[\begin{array}[]{cc}F_{o}\\ F_{s}\\ \end{array}\right]$ (1)
@@
3.2 Threat Model and Problem Formulation
@@
Problem. We assume the presence of a link recommender $f_{\phi}$ in the GNN, utilizing a subset of edges from core users in the original graph to suggest connections. This results in an updated graph $G_{r}$, with:
$G_{r}=(\hat{A},F_{p}),\quad\hat{A}=f_{\phi}(G_{p})$ (2)
@@
$\hat{A}=\left[\begin{array}[]{cc}A_{o}&A_{l}\\ {A}_{l}^{T}&A_{s}\\ \end{array}\right]$ (3)
@@ 3.3 Preliminary Analysis
$G_{s}=\text{Copy}(\mathcal{N}_{n}(v_{t})),\quad\mathcal{N}_{n}(v_{t})=\{u\in V|\text{dist}(u,v_{t})\leq n\}$ (5)
@@
Here, $\text{Copy}()$ is the replicating operation, $\mathcal{N}_{n}(v_{t})$ represent the n-hop neighborhoods set on $v_{t}$, and $\text{dist}(u,v_{t})$ gives the minimum distance between nodes $u$ and $v_{t}$. Such similarity seemingly facilitates the formation of links between the original graph and subgraph. However, simply generating links does not achieve our primary goal of reducing the accuracy of node classification. To address this, we introduce a perturbation in the form of Gaussian noise to the node features of the duplicated subgraph. By fine-tuning the number of hops and the intensity of perturbation, we can achieve a balance that facilitates link generation while reducing node classification accuracy, thereby achieving the adversarial objectives. Although we verified that GraphCopy can partially achieve the adversarial goal of SIA, it faces two significant limitations that reduce its effectiveness and practicality in real-world scenarios. First, simply duplicating the graph and altering node features does not ensure a high success rate for adversarial attacks. Second, replicating the n-hop neighborhood of a target node leads to a subgraph size that increases quadratically with the node degree. For nodes with high degrees, the attack budget increase substantially, making GraphCopy impractical for use in dense or large-scale networks. To maximize the effectiveness of attacks while minimizing costs, there is a strong need for an optimization-driven method to refine the subgraph for SIA.
4 Methodology
@@
In this section, we investigate an optimization-based approach for generating subgraphs to realize the SIA scenario. While extensive research has focused on targeting either node classification or link prediction models independently [13], the exploitation of the interaction between these two distinct model types to achieve an adversarial objective remains largely unexplored. Moreover, achieving dual adversarial goals within the SIA scenario presents significant technical challenges. Intuitively, generating links is more likely when there is a close resemblance in the node representations between the target victim nodes and the nodes of the aggressor subgraph, whereas a greater disparity in these representations increases the chances of the victim nodes being misclassified. As a result, a more deliberate subgraph optimization technique is essential to balance two conflicting objectives and identify the optimal subgraph for SIA. To this end, we introduce the LiSA framework, as depicted in Fig. 2.
Figure 2: An overview of LiSA framework.
4.1 LiSA Framework
@@
The LiSA framework is designed to challenge the performance of two GNN models through a strategic deployment of dual-surrogate models for bi-level optimization in a black-box attack context. We select the GAE [8] and GCN [9] as the surrogate models for link prediction and node classification, aiming to refine the adjacency matrix and feature attributes of the subgraph to meet two critical optimization goals concurrently: (i) the generation of adversarial edges between the original graph and the subgraph via the link recommender, and (ii) negatively affect node classification performance on target victim nodes once they are connected to the subgraph. To navigate these objectives, the attack loss function of LiSA includes two terms, balanced by a factor $\alpha$, to regulate the contribution of each model:
$\mathcal{L}_{atk}=\mathcal{L}^{cls}_{atk}(G_{r};\theta)+\alpha\mathcal{L}^{link}_{atk}(G_{p};\phi)$ (6)
@@
$\mathcal{L}^{link}_{atk}=\mathcal{L}_{recon}=-\sum_{i\in V}\sum_{j\in V}\log\hat{a}_{ij}-\sum_{n=1}^{Q}\cdot\mathrm{E}_{v_{n}\sim P_{n}(v_{i})}\log(1-\hat{{a_{n}}}_{ij})$ (7)
@@
Notably, the approach begins by pretraining a node classifier using a limited label set, subsequently swapping the top-2 output classes to define the attack labels [10]. This approach circumvents the traditional use of negative cross-entropy loss used in GradMax, which complicates the task of balancing the two objectives. The continuously increasing adversarial loss for classification, $\mathcal{L}^{cls}_{atk}$, usually fails to maintain balance between these two loss terms. By giving precedence to the second-highest output class for the attack, we foster a more stable training environment. The comprehensive training strategy is outlined as follows:
$\min_{A_{s},F_{s}}$ $\mathcal{L}^{cls}_{atk}(G_{r};\theta^{*})+\alpha\mathcal{L}^{link}_{atk}(G_{p};\phi^{*})$ (9)
s.t. $\theta^{*}=\mathop{\arg\min}_{\theta}\mathcal{L}^{cls}_{train}(f_{\theta}(G_{r})),\quad$ $\phi^{*}=\mathop{\arg\min}_{\phi}\mathcal{L}^{link}_{train}(f_{\phi}(G_{p})).$
@@
The method involves the simultaneous bi-level optimization of the two surrogate models, with optimizing space limited to the subgraph.
4.2 Training Algorithm